Grade 12Physical Chemistry

calculate the mole fraction of toulene in the vapour phase which is in equilib with a solution of benzene and tolene having a mole fraction of toulene 0.5.( the vapour pressure of pure benzene is 119 torr and that of toulene is 37 torr at the same temp).

Shriya Mehrotra
11 Years agoGrade 12

1 Answer

Naveen Kumar
11 Years ago
From Raoult’s law, the vapour pressure of components in a solution can be given as :
PA=P0* XA
where, A is an volatile component in the solution, X is the mole fraction of that component.
The total vapour pressure of the solution=Sum of vapour pressures due to all the components in the solution
Here P(total)=P(tolune) + P (Benzene)=P0* XT+P0* XB
Now frtom Dalton’s law of partial pressure, total pressure =sum of partial pressure of the components.and partial pressure=P(total)* Y
wehere, Y =mole fraction of a component in vapour phase.

So partial pressure of tolune=XT* PT0=YT*P(total) =YT* (PT+ PB)
Now mole fraction of tolune=0.5,and PT0=37 torr, we can also calculate PB=PB0*XB
So use the equation and calculate the result, But before that please try to understand this concept properly!!!!!